OpenStudy (anonymous):

Write an equation for the line that contains the point (3, 2) and is parallel to the line -4x – y = 6. A. -x + 4y = 5 B. 4x + y = -5 C. 4x + y = -1 D. 4x + y = 14

jimthompson5910 (jim_thompson5910):

anything parallel to -4x – y = 6 will be in the form -4x – y = C the left side stays the same while the right side will change

jimthompson5910 (jim_thompson5910):

-4x – y = C -4*3 – 2 = C ... plug in (x,y) = (3,2) simplify the left side to find C

OpenStudy (anonymous):

-14 ? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

yes, so -4x – y = C turns into -4x – y = -14

jimthompson5910 (jim_thompson5910):

-4x – y = -14 is a perfectly valid answer, but it's not listed but we can multiply both sides by -1 to get -4x – y = -14 -1*(-4x – y) = -1*(-14) 4x+y = 14

jimthompson5910 (jim_thompson5910):

so `-4x – y = -14` is equivalent to `4x+y = 14`

OpenStudy (anonymous):

so D and thank you @jim_thompson5910

jimthompson5910 (jim_thompson5910):

yep and yw